### v1

# Static Manifolds with Boundary and Rigidity of Scalar Curvature and Mean Curvature

Abstract.
On a compact manifold with boundary, the map consisting of the scalar curvature in the interior and the mean curvature on the boundary is a local surjection at generic metrics [C-V]. Moreover, this result may be localized to compact subdomains in an arbitrary Riemannian manifold with boundary [Sheng]. The non-generic case (also called non-generic domains) corresponds to static manifolds with boundary. We discuss their geometric properties, which also work as the necessary conditions of non-generic metrics. In space forms and the Schwarzschild manifold, we classify simple non-generic domains (with only one boundary component) and show their connection with rigidity theorems and the Schwarzschild photon sphere [Cederbaum, Cederbaum-G, Claudel-V-E].
1. Introduction
Let $(M^{n},g)$ be a complete Riemannian manifold. A static potential is a non-trivial solution $u\in C^{\infty}(M)$ to the equation
(1.1) $-\left(\Delta_{g}u\right)g+\operatorname{Hess}(u)-u\operatorname{Ric}_{g}=0.$
Riemannian manifolds admitting static potentials are called static manifolds. In fact, static potentials are related to static spacetimes in general relativity. A static spacetime is a four-dimensional Lorentzian manifold which possesses a timelike Killing field and a spacelike hypersurface which is orthogonal to the integral curves of this Killing field. Moreover, we have the following proposition:
Proposition 1.1 (Corvino [C]).
Let $(M^{n},g)$ be a complete Riemannian manifold. Then $u$ is a static potential if and only if the warped product metric $\bar{g}=-u^{2}\,dt^{2}+g$ is Einstein.
Static manifolds are interesting geometric objects that arise in other contexts as well. When Fischer and Marsden [F-M] studied the local surjectivity of the scalar curvature operator, they also derived the left hand side of equation (1.1) as the formal $L^{2}$-adjoint of the linearization of the scalar curvature. Later Corvino [C] considered localized deformations of metrics and extended the local surjectivity result at non-static metrics.
Static manifolds exhibit many interesting geometric properties. For example, its scalar curvature must be constant and the zero-set of a static potential is a totally geodesic hypersurface [C]. Therefore, many people are interested in the classification of static manifolds. In particular, Fischer and Marsden [F-M] raised the possibility of identifying all compact static manifolds. Now it is known that in dimension $3$, besides flat tori $T^{3}$ and round spheres $S^{3}$, $S^{1}\times S^{2}$ is also a compact static manifold. Later, other warped metrics on $S^{1}\times_{r}S^{2}$ were found to be static [Ejiri]. Kobayashi [K] and Lafontaine [Lafontaine] proved a more general classification result for locally conformally flat static manifolds. Qing and Yuan [Q-Y] gave a classification result in dimension $3$ assuming a weaker hypothesis on the Cotton tensor. On the other hand, Ambrozio [Ambrozio] computed a Bochner type formula and deduced classification results for compact static three-manifolds with positive scalar curvature.
It is natural to extend the above discussion to manifolds with boundary. Of course there are different ways to define so-called “static manifolds with boundary” with different boundary conditions. In [Sheng], the author followed the ideas of Escobar [E1, E2] and Corvino [C] to consider localized deformations prescribing scalar curvature and boundary mean curvature. It turns out that the map is a local surjection at generic metrics. So we are interested to know the properties of non-generic metrics when the map is not a local surjection. To be more precise, let $(\Omega,\Sigma)$ be a domain in $(M^{n},g)\,(n\geq 3)$ where $\Sigma$ is the boundary of $\Omega$. Then $(\Omega,\Sigma)$ is called a non-generic domain (or “static manifolds with boundary” in our setting) if there is a non-trivial solution $u\in C^{\infty}(M)$ to the boundary value problem:
(1.2) $-\left(\Delta_{g}u\right)g+\operatorname{Hess}(u)-u\operatorname{Ric}_{g}&=0\qquad&\text{in }\Omega\\ u_{\nu}\hat{g}&=uh&\text{on }\Sigma.\end{aligned}\right.$
Here $\hat{g}$ is the induced metric on $\Sigma$, $h$ is the second fundamental form of $\Sigma$, and $\nu$ is the outward unit normal. Note that the interior equation is the same as (1.1), while the boundary equation is closely related to the linearization of the mean curvature operator (see [Sheng] for more details). Henceforth, we may interchangeably use “static manifolds with boundary” and “non-generic domains” to denote the same concept.
With the above boundary condition, static manifolds with boundary have more interesting properties and geometric structures than the closed case. So in Section 2 and 3, we will discuss some global results and mainly focus on compact static manifolds with boundary.
It is worth noting that static manifolds with boundary may not be compact in general. And it turns out that non-compact static manifolds with boundary are often related to the rigidity theorems of certain kinds of positive mass theorems [A-B-dL, A-dL, H-M-R, M, S-T, H-W, Q-Y2]. Moreover, they tend to have special meanings in general relativity. Therefore, we will try to classify non-generic domains in space forms and in the Schwarzschild manifold, and show their connection with rigidity theorems and the Schwarzschild photon sphere [Claudel-V-E, Cederbaum, Cederbaum-G].
The paper is organized as follows. In Sect. 2, we define basic notation and review previous results about non-generic domains. In Sect. 3, we focus on compact non-generic domains and discuss global properties. In Sect. 4, we use conformal methods to classify simple non-generic domains (with only one boundary component) in space forms and in the Schwarzschild manifold. We then discuss how to construct general non-generic domains with multiple boundary components.
2. Preliminaries
In this section, we will review some basic properties of non-generic domains.
Let $(\Omega,\Sigma)$ be a domain in $(M^{n},g)\,(n\geq 3)$ where $\Sigma$ is the boundary of $\Omega$. Then $(\Omega,\Sigma)$ is called a non-generic domain if there is a non-trivial solution $u\in C^{\infty}(M)$ to the boundary value problem (1.2). Many authors call a non-trivial solution to (1.1) a static potential. With the boundary condition, however, we will call a non-trivial solution to (1.1) a possible static potential on $(\Omega,\Sigma)$, and if it satisfies the boundary condition on $\Sigma$ as well, we then call it a static potential on $(\Omega,\Sigma)$ (see [A-dL, A-dL2] for similar definitions of static potentials and static manifolds with boundary). That is, a non-trivial solution to (1.2) is called a static potential in our setting. So in this case, generic conditions [Sheng] are not satisfied.
To make the notation consistent, recall in [Sheng] we defined the operator $L^{*}:C^{\infty}(\Omega\cup\Sigma)\longrightarrow\tilde{\mathcal{S}}^{(0,2)}(\Omega\cup\Sigma)$ by
$L^{*}u=-\left(\Delta_{g}u\right)g+\operatorname{Hess}(u)-u\operatorname{Ric}_{g},$
and the operator $\Phi^{*}:C^{\infty}(\Omega\cup\Sigma)\longrightarrow\tilde{\mathcal{S}}^{(0,2)}(\Omega\cup\Sigma)\times\tilde{\mathcal{S}}^{(0,2)}(\Sigma)$ by
$\Phi^{*}u=(L^{*}u,u_{\nu}\hat{g}-uh).$
Here $\tilde{\mathcal{S}}^{(0,2)}(\Omega)$ denotes the space of symmetric $(0,2)$-tensor fields on $\Omega$ that are smooth. With these notation, the space of static potentials on $(\Omega,\Sigma)$ can now be denoted as $\operatorname{ker}\Phi^{*}$, and $(\Omega,\Sigma)$ is a non-generic domain if and only if $\operatorname{ker}\Phi^{*}\neq 0$.
Let us first recall some geometric properties of non-generic domains derived in [Sheng].
Theorem 2.1 ([Sheng]).
If $(\Omega,\Sigma)$ is a non-generic domain, then the scalar curvature of $\Omega$ is constant, the boundary $\Sigma$ is umbilic, and the mean curvature is constant on $\Sigma$.
Remark 2.2.
The same proof in [Sheng] actually gives $\operatorname{Ric}_{i\nu}=0\,\,\,(i=1,\ldots,n-1)$ on $\Sigma$ as well.
Proposition 2.3 ([Sheng]).
For non-generic domains $(\Omega,\Sigma)$ in $(M^{n},g)$, we have $\operatorname{dim}\operatorname{ker}\Phi^{*}\leq n$ in $H_{\mathrm{loc}}^{2}(\Omega)$.
As we will see in section 4, the upper bound is achieved by simple non-generic domains in space forms.
Proposition 2.4 ([Sheng]).
$H_{\mathrm{loc}}^{2}$ elements in $\operatorname{ker}\Phi^{*}$ are actually in $C^{2}(\bar{\Omega})$.
In [Sheng] we also derived a very useful lemma.
Lemma 2.5 ([Sheng]).
There is no non-trivial solution in $H_{\operatorname{loc}}^{2}(\Omega)$ to the following equations:
{$L^{*}u&=0\qquad\,\text{in }\Omega\\ u=u_{\nu}&=0\qquad\text{on }\Sigma.\end{aligned}\right.$
As a quick corollary, for non-generic domains with totally geodesic boundary, the space of static potentials is fully determined by the boundary data on $\Sigma$.
Corollary 2.6.
Let $(\Omega,\Sigma)$ be a non-generic domain. If $H=0$, then $\Sigma$ is totally geodesic, and $\operatorname{ker}\Phi^{*}$ is fully determined by the boundary data on $\Sigma$.
Proof.
If $H=0$, then for any $u\in\operatorname{ker}\Phi^{*}$, $L^{*}u=0$ in $\Omega$ and $u_{\nu}=0$ on $\Sigma$. So for any $u_{1},u_{2}\in\operatorname{ker}\Phi^{*}$ with $u_{1}\equiv u_{2}$ on $\Sigma$, consider $v=u_{1}-u_{2}\in\operatorname{ker}\Phi^{*}$. We have
{$L^{*}v&=0\qquad\,\text{in }\Omega\\ v=v_{\nu}&=0\qquad\text{on }\Sigma.\end{aligned}\right.$
By Lemma 2.5, $v\equiv 0$ in $\Omega$. This means $u_{1}\equiv u_{2}$ in $\Omega$. ∎
3. Compact non-generic domains
Let us now focus on compact non-generic domains.
Taking the trace of (1.2), we get
(3.1) $\Delta_{g}u+\frac{R_{g}}{n-1}u&=0\quad&\text{in }\Omega\\ u_{\nu}-\frac{H}{n-1}u&=0&\text{on }\Sigma.\end{aligned}\right.$
This means, for any static potential $u$, they must first satisfy (3.1). It is very useful when we want to get some global properties of compact non-generic domains.
From Theorem 2.1 we know that the scalar curvature $R$ of $\Omega$ and the mean curvature $H$ of $\Sigma$ are constant on non-generic domains. Thus all the non-generic domains may be divided into 9 cases, according to whether $R$ and $H$ are positive, negative or zero. However, some of the cases are not possible. Cruz and Vitório [C-V] found that:
Proposition 3.1 (Cruz-Vitório).
Let $(\Omega,\Sigma)$ be a compact non-generic domain.
(i) If $R=0$, then $H\geq 0$; and if $H=0$ as well, then $h\equiv 0$.
(ii) If $H=0$, then $R\geq 0$; and if $R=0$ as well, then $Ric_{g}\equiv 0$.
Note that:
(i) If $R=0$ and $H>0$, then $\frac{H}{n-1}$ is an Steklov eigenvalue; in this case, $\operatorname{ker}\Phi^{*}$ lies in the eigenspace corresponding to the Steklov eigenvalue $\frac{H}{n-1}$.
(ii) If $R>0$ and $H=0$, then $\frac{R}{n-1}$ is an eigenvalue of the Neumann boundary value problem; in this case, $\operatorname{ker}\Phi^{*}$ lies in the eigenspace corresponding to the eigenvalue $\frac{R}{n-1}$.
(iii) If $R=H=0$, then $g$ is Ricci flat and $\Sigma$ is totally geodesic. On the other hand, Ho and Huang [H-H] found that:
Proposition 3.2 (Ho-Huang).
If $g$ is Ricci flat and $\Sigma$ is totally geodesic, then $\operatorname{ker}\Phi^{*}\neq 0$. In fact, $\operatorname{ker}\Phi^{*}$ consists of constant functions in $\Omega$.
Thus for case (iii) we have the following result:
Theorem 3.3.
Let $(\Omega,\Sigma)$ be a compact non-generic domain. If $R=H=0$, then $g$ is Ricci flat, $\Sigma$ is totally geodesic, and $\operatorname{ker}\Phi^{*}$ consists of constant functions in $\Omega$.
In this case, we actually have the boundary scalar curvature is also 0.
Corollary 3.4.
Let $(\Omega,\Sigma)$ be a compact non-generic domain. If $R=H=0$, then the boundary scalar curvature $R_{\Sigma}=0$.
Proof.
By looking at the $\nu\nu$-th term of $L^{*}u=0$, we have
$\Delta_{g}u+u_{\nu\nu}-uR_{\nu\nu}=0.$
Since $u$ is a non-zero constant, $R_{\nu\nu}$ must be 0. Then by the Gauss equation,
$R_{\Sigma}=R-2R_{\nu\nu}-\|h\|^{2}+H^{2}=0.$
We may further rule out some more cases:
Theorem 3.5.
Let $(\Omega,\Sigma)$ be a compact non-generic domain. If $R<0$, then $H>0$; if $H<0$, then $R>0$.
Proof.
If $0\neq u\in\operatorname{ker}\Phi^{*}$, then $u$ must satisfy (3.1). Multiplying $u$ to the first equation and integrating it over $\Omega$, we get
[…]
Hence, if $R<0$, the first integral is strictly negative, then $H$ must be strictly positive. Similarly, if $H<0$, then $R>0$. ∎
As a summary, we have the following diagram for compact non-generic domains:
H R $+$ 00 --
$+$ ? Steklov ?
00 Neumann * $\times$
-- ? $\times$ $\times$
Table 1. Cases of compact non-generic domains
In the next section, we will see some more examples that will fill up the diagram above.
4. General non-generic domains
Now we would like to study non-generic domains which may not be compact. Among them we are particularly interested in non-generic domains in a space form and in the Schwarzschild manifold.
The umbilic hypersurface in a space form has constant mean curvature [dC], thus the region bounded by an umbilic hypersurface in a space form satisfies all the necessary conditions of a non-generic domain. We will see that the regions bounded by umbilic hypersurfaces in a space form are indeed non-generic domains.
For simplicity, we will first consider non-generic domains with only one boundary component. We call them simple non-generic domains. If a non-generic domain has multiple boundary components, then it can be viewed as the intersection of simple non-generic domains, as long as their boundaries do not intersect with each other.
4.1. The unit sphere $\mathbb{S}^{n}$
Consider the standard metric $g_{\mathbb{S}^{n}}$ on the unit sphere $\mathbb{S}^{n}$. In this case, the sectional curvature $K=1$, the Ricci curvature $\operatorname{Ric}=(n-1)g_{\mathbb{S}^{n}}$, and the scalar curvature $R=n(n-1)$.
Cruz and Vitório [C-V] found that the upper hemisphere with its boundary is a non-generic domain in $\mathbb{S}^{n}$. Later Ho and Huang [H-H] had the following result:
Proposition 4.1 (Ho-Huang).
Let $(\mathbb{S}^{n}_{+},\partial\mathbb{S}^{n}_{+})$ be the $n$-dimensional upper hemisphere equipped with the standard metric $g_{\mathbb{S}^{n}}$, where $n\geq 3$. Then $\operatorname{ker}\Phi^{*}$ is non-trivial. Moreover,
$\operatorname{ker}\Phi^{*}=\operatorname{span}\{x_{1},\cdots,x_{n}\},$
where $(x_{1},\cdots,x_{n+1})$ are the coordinates of $\mathbb{S}^{n}\subset\mathbb{R}^{n+1}$.
In fact, for $x_{i}\,(i=1,\cdots,n+1)$, we have
$\operatorname{Hess}_{g_{\mathbb{S}^{n}}}x_{i}+x_{i}g_{\mathbb{S}^{n}}=0.$
Thus,
$L^{*}(x_{i})$ $=-\left(\Delta x_{i}\right)g_{\mathbb{S}^{n}}+\operatorname{Hess}_{g_{\mathbb{S}^{n}}}x_{i}-(n-1)x_{i}g_{g_{\mathbb{S}^{n}}}$
$=-\left(\Delta x_{i}\right)g_{\mathbb{S}^{n}}-x_{i}g_{\mathbb{S}^{n}}-(n-1)x_{i}g_{g_{\mathbb{S}^{n}}}$
$=-\left(\Delta x_{i}+nx_{i}\right)g_{\mathbb{S}^{n}}$
$=0.$
Proposition 4.2.
The linear combinations of the coordinates $x_{i}\,(i=1,\cdots,n+1)$ are the only possible static potentials on non-generic domains $\Omega\subset\mathbb{S}^{n}\,(n\geq 3)$.
Proof.
From the non-generic condition on $\Omega$
$L^{*}u=-\left(\Delta u\right)g+\operatorname{Hess}u-(n-1)ug=0,$
we get
$\mathcal{L}_{Du}g=2\operatorname{Hess}u=2\left(\Delta u+(n-1)u\right)g.$
Therefore, $Du$ is a conformal killing vector field.
From Liouville’s theorem on conformal mappings [B-P], every conformal diffeomorphism between two domains of $\mathbb{R}^{n}\,(n\geq 3)$ has the form
$\mathbf{x}\mapsto\lambda Ai(\mathbf{x})+\mathbf{b}$
where $\lambda>0,A\in O(n)$, $i$ is either the identity or an inversion, and $\mathbf{b}\in\mathbb{R}^{n}$.
Moreover, for conformal groups of $\mathbb{S}^{n}\,(n\geq 3)$, we have [L-Y]
$\operatorname{Conf}(\mathbb{S}^{n})/SO(n+1,1)\cong\operatorname{B}^{n+1},$
and similar results hold for open subsets of $\mathbb{S}^{n}$.
However, a smooth gradient flow cannot be a family of inversions, and it cannot be a family of rotations either. This is because, if $Du$ is a generator in $\operatorname{Isom}(\mathbb{S}^{n})$, then
$0=\mathcal{L}_{Du}g=2\operatorname{Hess}u,$
but
$2\left(\Delta u+(n-1)u\right)g=0+2(n-1)ug\neq 0,$
which contradicts the equation of the non-generic condition.
Thus $Du$ can only lie in the Lie algebra of $F(\operatorname{B}^{n+1})\subset\operatorname{Conf}(\mathbb{S}^{n})$[L-Y]. Any point $0\neq\mathbf{a}\in\operatorname{B}^{n+1}$ can be written as $\mathbf{a}=c_{1}e_{1}+\cdots+c_{n+1}e_{n+1}$, and the associated conformal vector field is given by $V_{a}=\frac{1}{\|\mathbf{a}\|}(c_{1}e_{1}^{\top}+\cdots+c_{n+1}e_{n+1}^{\top})$. Here $\{e_{1},\cdots,e_{n+1}\}$ is an orthonormal basis of $\mathbb{R}^{n+1}$, and $e_{i}^{\top}$’s are the projections onto $T\mathbb{S}^{n}$. This means,
$Du=c_{1}e_{1}^{\top}+\cdots+c_{n+1}e_{n+1}^{\top},$
which implies
$u=c_{1}x_{1}+\cdots+c_{n+1}x_{n+1}.$
As a conclusion, the static potential $u$ can only be a linear combination of the coordinate functions $x_{i}$’s. ∎
On the other hand, it is a well-known fact that a compact (without boundary) totally umbilical hypersurface of a simply connected real space form is a geodesic sphere. So the only possible non-generic domains in $\mathbb{S}^{n}$ are spherical caps.
Theorem 4.3.
Let $(\Omega,\Sigma)$ be the $n$-dimensional spherical cap equipped with the standard metric $g_{\mathbb{S}^{n}}$, where $n\geq 3$. Then $\operatorname{ker}\Phi^{*}$ is non-trivial. Moreover, if the north pole $N$ is the center of the spherical cap, then
$\operatorname{ker}\Phi^{*}=\operatorname{span}\{x_{1},\cdots,x_{n}\}.$
In fact, the spherical caps are the only simple non-generic domains in $\mathbb{S}^{n}$.
Proof.
As we discussed above, the coordinates $x_{i}\,(i=1,\cdots,n)$ are the only possible static potentials, and they satisfy (1.1).
Let us now check $x_{i}$’s also satisfy the boundary equation of (1.2). If the north pole $N$ is the center of the spherical cap, then $\Sigma$ may be parametrized as $\{\mathbf{x}\in\mathbb{S}^{n}:x_{n+1}=\operatorname{cos}\theta\}$, so that any point $\mathbf{x}\in\Sigma$ can be written as $\mathbf{x}=(\operatorname{sin}\theta\,\mathbf{\xi},\operatorname{cos}\theta)$, where $\mathbf{\xi}\in\mathbb{S}^{n-1}$, and the induced metric on $\Sigma$ is $\hat{g}=\operatorname{sin}^{2}\theta\,g_{\mathbb{S}^{n-1}}$.
In this case, the outward unit normal $\nu=\frac{\partial}{\partial\theta}$, and the mean curvature may be calculated as:
$H$ $=\frac{\partial}{\partial\theta}\operatorname{ln}\sqrt{\operatorname{det}\hat{g}}$
$=\frac{\partial}{\partial\theta}\operatorname{ln}\sqrt{\left(\operatorname{sin}^{2}\theta\right)^{n-1}\operatorname{det}g_{\mathbb{S}^{n-1}}}$
$=\frac{\partial}{\partial\theta}\left((n-1)\operatorname{ln}(\operatorname{sin}\theta)+\operatorname{ln}\sqrt{\operatorname{det}g_{\mathbb{S}^{n-1}}}\right)$
$=(n-1)\frac{\operatorname{cos}\theta}{\operatorname{sin}\theta}$
This means,
$(x_{i})_{\nu}-\frac{H}{n-1}x_{i}=\operatorname{cos}\theta\,\xi_{i}-\frac{\operatorname{cos}\theta}{\operatorname{sin}\theta}\operatorname{sin}\theta\,\xi_{i}=0\qquad\text{on }\Sigma$
Thus,
$\operatorname{ker}\Phi^{*}=\operatorname{span}\{x_{1},\cdots,x_{n}\}.$
For spherical caps, the scalar curvature $R>0$; when $\theta$ varies between 00 and $\pi$, the range of the mean curvature $H$ is the set of all real numbers $\mathbb{R}$.
We would like to mention here that non-generic domains in $\mathbb{S}^{n}$ are related to the famous Min-Oo’s Conjecture [M-O]:
Conjecture 4.4 (Min-Oo).
Let $(M,g)$ be an $n$-dimensional compact Riemannian manifold with boundary and the scalar curvature $R\geq n(n-1)$. The boundary is isometric to the standard sphere $\mathbb{S}^{n-1}$ and is totally geodesic. Then $(M,g)$ is isometric to the hemisphere $\mathbb{S}^{n}_{+}$.
Min-Oo’s conjecture has been verified in many special cases (see e.g. [Brendle-M, E, H-W, H-W2, M-T]), however, counterexamples were also constructed [Brendle-M-N, C-E-M].
4.2. The Euclidean space $\mathbb{R}^{n}$
Consider the flat metric $g_{ij}=\delta_{ij}$ on the Euclidean space $\mathbb{R}^{n}$.
Cruz and Vitório [C-V] found that the Euclidean ball of radius $r$ with its boundary is a non-generic domain in $\mathbb{R}^{n}$. Later Ho and Huang [H-H] had the following result:
Proposition 4.5 (Ho-Huang).
Let $(\operatorname{B}^{n},\partial\operatorname{B}^{n})$ be the $n$-dimensional unit ball equipped with the flat metric. Then $\operatorname{ker}\Phi^{*}$ is non-trivial. Moreover,
$\operatorname{ker}\Phi^{*}=\operatorname{span}\{x_{1},\cdots,x_{n}\},$
where $(x_{1},\cdots,x_{n})$ are the coordinates of $\operatorname{B}^{n}\subset\mathbb{R}^{n}$.
In fact, the same proof as in [H-H] would tell us that, if $(\Omega,\Sigma)=(B_{r},S_{r})$ is the Euclidean ball of radius $r$ with its boundary, then
$\operatorname{ker}\Phi^{*}=\operatorname{span}\{x_{1},\cdots,x_{n}\}.$
Thus, the interior region of a Euclidean sphere is a non-generic domain in $\mathbb{R}^{n}$.
Let us try to find other non-generic domains in $\mathbb{R}^{n}$. First of all, notice that we have the following result:
Proposition 4.6.
The linear combinations of the coordinates $x_{i}\,(i=1,\cdots,n)$ and constant functions are the only possible static potentials on non-generic domains $\Omega\subset\mathbb{R}^{n}$.
Proof.
The coordinates $x_{i}$ as well as constant functions are possible static potentials, because
$L^{*}(x_{i})=-\left(\Delta x_{i}\right)g+\operatorname{Hess}x_{i}-x_{i}\operatorname{Ric}_{g}=0,$
$\text{and}\quad L^{*}(c)=-\left(\Delta c\right)g+\operatorname{Hess}c-c\operatorname{Ric}_{g}=0.$
On the other hand, they are the only possible static potentials. Since for any static potential $u$,
$L^{*}u=-\left(\Delta u\right)g+\operatorname{Hess}u=0.$
From (3.1) we know $\Delta u=0$. So this means $\operatorname{Hess}u=0$, which implies $u$ is a linear combination of the coordinates $x_{i}$ and constant functions. ∎
As for totally umbilical hypersurfaces $\Sigma\subset\mathbb{R}^{n}$, they can only be spheres or planes [dC]. So let us check that the exterior region of a Euclidean sphere and the upper half plane (with their boundary) are also non-generic domains in $\mathbb{R}^{n}$.
If $\Omega$ is the exterior region of a sphere of radius $r$ and $\Sigma=S_{r}$, then
$(x_{i})_{\nu}-\frac{H}{n-1}x_{i}=Dx_{i}\cdot\left(-\frac{x}{r}\right)+\frac{1}{r}x_{i}=-\frac{x_{i}}{r}+\frac{x_{i}}{r}=0\qquad\text{on }\Sigma$
Thus, the exterior region of a sphere is a non-generic domain in $\mathbb{R}^{n}$, and
$\operatorname{ker}\Phi^{*}=\operatorname{span}\{x_{1},\cdots,x_{n}\}.$
If $\Omega$ is the upper half space $\{x_{n}>0\}$ and $\Sigma=\{x_{n}=0\}$, then for $11\leq i\leq n-$ 1,
$(x_{i})_{\nu}-\frac{H}{n-1}x_{i}=\frac{\partial x_{i}}{\partial x_{n}}+0=0\qquad\text{on }\Sigma$
$\text{and}\quad(c)_{\nu}-\frac{H}{n-1}c=0+0=0\qquad\text{on }\Sigma$
Thus, the upper half space is a non-generic domain in $\mathbb{R}^{n}$, and
$\operatorname{ker}\Phi^{*}=\operatorname{span}\{1,x_{1},\cdots,x_{n-1}\}.$
As a summary, we have the following classification result:
Theorem 4.7 (Classification of simple non-generic domains in $\mathbb{R}^{n}$).
There are two types of boundaries of non-generic domains in $\mathbb{R}^{n}$:
(i) the sphere;   (ii) the hyperplane.
In 2016, Almaraz-Barbosa-de Lima [A-B-dL] had the following conjecture of a positive mass theorem for asymptotically flat manifolds with a non-compact boundary, and they were able to prove the case when either $3\leq n\leq 7$ or $n\geq 3$ and $M$ is spin.
Conjecture 4.8 (Almaraz-Barbosa-de Lima).
If $(M,g)$ is asymptotically flat and satisfies $R_{g}\geq 0$ and $H_{g}\geq 0$, then $m(M,g)\geq 0$, with the equality occurring if and only if $(M,g)$ is isometric to (+$(\mathbb{R}^{n}_{+},\delta)$.
They also had the following proposition regarding the rigidity part:
Proposition 4.9 (Almaraz-Barbosa-de Lima).
Let $(M,g)$ be asymptotically flat with $3\leq n\leq 7$ or $n\geq 3$ and $M$ spin, and satisfy $R_{g}\geq 0$ and $H_{g}\geq 0$. Assume further that there exists a compact subset $K\subset M$ such that $(M\setminus K,g)$ is isometric to (+$(\mathbb{R}^{n}_{+}\setminus\bar{B}_{1}^{+}(0),\delta)$. Then $(M,g)$ is isometric to (+$(\mathbb{R}^{n}_{+},\delta)$.
Thanks to the above proposition, we can now better understand why the half space is a non-generic domain in $\mathbb{R}^{n}$. Suppose not, then the half space with $R=0$ and $H=0$ satisfies generic conditions. This means we would be able to perform a localized deformation [Sheng] on a compact region $K\subset\mathbb{R}^{n}_{+}$ such that $R>0$ or $H>0$ in $K$. However, this is a contradiction to the rigidity theorem.
There is another generalized version of the positive mass theorem [M, S-T] that will imply the following rigidity theorem for the unit ball in $\mathbb{R}^{n}$:
Proposition 4.10 (Hang-Wang [H-W]).
Let $(M,g)$ be an $n$-dimensional compact Riemannian manifold with boundary and the scalar curvature $R\geq 0$. The boundary is isometric to the standard sphere $\mathbb{S}^{n-1}$ and has mean curvature $n-1$. Then $(M,g)$ is isometric to the unit ball in $\mathbb{R}^{n}$. (If $n>7$, we also assume $M$ is spin.)
Notice that it can also give a good explanation of why the unit ball is a non-generic domain in $\mathbb{R}^{n}$.
4.3. The hyperbolic space $\mathbb{H}^{n}$
Consider the hyperboloid model for the hyperbolic space $\mathbb{H}^{n}$. To be more precise, let us consider the Minkowski quadratic form defined on $\mathbb{R}^{1,n}$ by
$Q\left(x_{0},x_{1},\ldots,x_{n}\right)=-x_{0}^{2}+x_{1}^{2}+\cdots+x_{n}^{2}.$
Then the hyperbolic space is defined by
$\mathbb{H}^{n}=\{\mathbf{x}\in\mathbb{R}^{1,n}:Q(\mathbf{x})=-1,x_{0}>0\}$
with induced metric $g_{\mathbb{H}^{n}}$. In this case, the sectional curvature $K=-1$, the Ricci curvature $\operatorname{Ric}=-(n-1)g_{\mathbb{H}^{n}}$, and the scalar curvature $R=-n(n-1)$.
As a part of the unit sphere in the Minkowski space, the hyperbolic space has many properties that are similar to the unit sphere $\mathbb{S}^{n}$ in the Euclidean space.
Proposition 4.11.
The linear combinations of the coordinates $x_{i}\,(i=0,1,\cdots,n)$ are the only possible static potentials on non-generic domains $\Omega\subset\mathbb{H}^{n}$. Here $(x_{0},x_{1},\cdots,x_{n})$ are the coordinates of $\mathbb{H}^{n}\subset\mathbb{R}^{1,n}$.
Proof.
Let us consider the hyperbolic space as a graph in $\mathbb{R}^{1,n}$. To be more precise, any point $(x_{0},\mathbf{x})=(x_{0},x_{1},\ldots,x_{n})\in\mathbb{H}^{n}$ satisfies
$x_{0}=\sqrt{1+\|\mathbf{x}\|^{2}}$
Thus the hyperbolic space can be characterized as
$F:\mathbb{R}^{n}$ $\to\mathbb{R}^{1,n};$
$\mathbf{x}$ $\mapsto(\sqrt{1+\|\mathbf{x}\|^{2}},\mathbf{x})$
Denote $e_{i}\in T\mathbb{H}^{n}\,(i=1,\cdots,n)$ as
$e_{i}=dF\left(\frac{\partial}{\partial x_{i}}\right)=\left(\frac{x_{i}}{\sqrt{1+\|\mathbf{x}\|^{2}}},0,\cdots,\mathop{1}\limits_{i\text{-th}},\cdots,0\right)=\frac{x_{i}}{x_{0}}\frac{\partial}{\partial x_{0}}+\frac{\partial}{\partial x_{i}}.$
Then the induced metric is
$g_{ij}=\left<e_{i},e_{j}\right>=\delta_{ij}-\frac{x_{i}x_{j}}{x_{0}^{2}}\qquad\text{and}\qquad g^{ij}=\delta^{ij}+x_{i}x_{j}.$
For $k=1,\cdots,n$
$\operatorname{Hess}_{ij}x_{k}$ $=e_{i}(e_{j}x_{k})-\Gamma_{ij}^{l}e_{l}x_{k}$
$=e_{i}(\delta_{jk})-\Gamma_{ij}^{l}\delta_{lk}$
$=-\Gamma_{ij}^{k}$
$=-\frac{1}{2}g^{kl}\left(\frac{\partial}{\partial x^{j}}g_{li}+\frac{\partial}{\partial x^{i}}g_{lj}-\frac{\partial}{\partial x^{l}}g_{ij}\right)$
$\frac{\partial}{\partial x^{j}}g_{li}$ $=\frac{\partial}{\partial x^{j}}\left(\delta_{li}-\frac{x_{l}x_{i}}{x_{0}^{2}}\right)$
$=-\frac{\frac{\partial}{\partial x^{j}}(x_{l}x_{i})x_{0}^{2}-2x_{l}x_{i}x_{0}\frac{\partial}{\partial x^{j}}x_{0}}{x_{0}^{4}}$
$=-\frac{\delta_{jl}x_{i}+\delta_{ji}x_{l}}{x_{0}^{2}}+2\frac{x_{l}x_{i}x_{j}}{x_{0}^{4}}$
So
$\operatorname{Hess}_{ij}x_{k}$ $=\frac{1}{2}g^{kl}\left(\frac{\delta_{jl}x_{i}+\delta_{ji}x_{l}}{x_{0}^{2}}-2\frac{x_{l}x_{i}x_{j}}{x_{0}^{4}}+\frac{\delta_{il}x_{j}+\delta_{ji}x_{l}}{x_{0}^{2}}-2\frac{x_{l}x_{i}x_{j}}{x_{0}^{4}}\right)$
$\quad+\frac{1}{2}g^{kl}\left(-\frac{\delta_{jl}x_{i}+\delta_{li}x_{j}}{x_{0}^{2}}+2\frac{x_{l}x_{i}x_{j}}{x_{0}^{4}}\right)$
$=(\delta^{kl}+x^{k}x^{l})\left(\frac{\delta_{ji}x_{l}}{x_{0}^{2}}-\frac{x_{l}x_{i}x_{j}}{x_{0}^{4}}\right)$
$=\frac{\delta_{ji}x_{k}}{x_{0}^{2}}+\frac{\sum_{l}x_{k}\delta_{ji}x_{l}^{2}}{x_{0}^{2}}-\frac{x_{k}x_{i}x_{j}}{x_{0}^{4}}-\frac{\sum_{l}x_{k}x_{i}x_{j}x^{2}_{l}}{x_{0}^{4}}$
$=x_{k}\delta_{ij}-\frac{x_{k}x_{i}x_{j}}{x_{0}^{2}}$
$=x_{k}g_{ij}$
and
$\operatorname{Hess}_{ij}x_{0}$ $=e_{i}(e_{j}x_{0})-\Gamma_{ij}^{l}e_{l}x_{0}$
$=e_{i}\left(\frac{x_{j}}{x_{0}}\right)-\Gamma_{ij}^{l}\frac{x_{l}}{x_{0}}$
$=\frac{\delta_{ij}x_{0}-x_{j}\frac{x_{i}}{x_{0}}}{x_{0}^{2}}+\sum_{l}x_{l}g_{ij}\frac{x_{l}}{x_{0}}$
$=\frac{1}{x_{0}}g_{ij}+\frac{1}{x_{0}}g_{ij}\sum_{l}x^{2}_{l}$
$=x_{0}g_{ij}$
Thus for $i=0,1,\cdots,n$
$\operatorname{Hess}_{g_{\mathbb{H}^{n}}}x_{i}=x_{i}g_{\mathbb{H}^{n}},$
and we have
$L^{*}(x_{i})=-\left(\Delta x_{i}\right)g+x_{i}g+(n-1)x_{i}g=\left(-\Delta x_{i}+nx_{i}\right)g=0.$
On the other hand, by Corvino [C]*Corollary 2.4 we have $\operatorname{dim}\operatorname{ker}L^{*}\leq n+1$ in $H_{\mathrm{loc}}^{2}(\Omega)$. This means the linear combinations of the coordinates $x_{i}\,(i=0,1,\cdots,n)$ are the only possible static potentials on $\Omega$. ∎
As for totally umbilical hypersurfaces $\Sigma\subset\mathbb{H}^{n}$, they are of the form $P\cap\mathbb{H}^{n}$, where $P$ is some affine hyperplane of $\mathbb{R}^{1,n}$ [dC]. Alternatively, if we consider the upper half-space model, then $\Sigma$ can only be the geodesic spheres, the horospheres, the hyperspheres and the intersection with $\mathbb{H}^{n}$ of hyperplanes of $\mathbb{R}^{n}$ [dC]. To be more precise, following do Carmo, let $\Sigma=S\cap\mathbb{H}^{n}$ be the intersection of $\mathbb{H}^{n}$ with a Euclidean $(n-1)$-sphere $S\subset\mathbb{R}^{n}$ of radius $1$ and center in $\mathbb{H}^{n}$. Then the mean curvature $\frac{H}{n-1}$ of $\Sigma$ is
(i) $1$, if $S$ is tangent to $\partial\mathbb{H}^{n}$ (horosphere);
(ii) $\operatorname{cos}\alpha$, if $S$ makes an angle $\alpha$ with $\partial\mathbb{H}^{n}$ (hypersphere);
(iii) the height $h$ of the Euclidean center of $S$ relative to $\partial\mathbb{H}^{n}$, if $S\subset\mathbb{H}^{n}$ (geodesic sphere).
$\alpha$$O$$\partial\mathbb{H}^{n}$
Figure 1. A hypersphere in the half-space model
Now let us check one by one whether each of the above umbilical hypersurfaces will give us a boundary of the non-generic domain in $\mathbb{H}^{n}$. However, Proposition 4.11 tells us that the coordinates of $\mathbb{H}^{n}\subset\mathbb{R}^{1,n}$ in the hyperboloid model are the only possible static potentials, so we would like to first rewrite these coordinates $x_{i}$’s in terms of the coordinates in the upper half-space model.
Denote $\mathbb{H}^{n}_{-1}$ as the hyperboloid model, $D^{n}$ as the unit ball model, and $\mathbb{H}^{n}$ as the upper half-space model of the hyperbolic space. The inverse of the stereographic projection
$f:D^{n}$ $\to\mathbb{H}^{n}_{-1};$
$(0,u_{1},\cdots,u_{n})$ $\mapsto(x_{0},x_{1},\cdots,x_{n})$
is given by
$x_{\alpha}$ $=\frac{2u_{\alpha}}{1-\sum_{\alpha}u_{\alpha}^{2}},\quad\alpha=1,\cdots,n$
$x_{0}$ $=\frac{2}{1-\sum_{\alpha}u_{\alpha}^{2}}-1.$
And the conformal mapping between the unit ball $D^{n}$ and the upper half-space $\mathbb{H}^{n}$
$g:\mathbb{H}^{n}$ $\to D^{n};$
$(y_{1},\cdots,y_{n})$ $\mapsto(u_{1},\cdots,u_{n})$
is given by
$u_{k}$ $=\frac{2y_{k}}{\sum_{k}y_{k}^{2}+(y_{n}+1)^{2}},\qquad k=1,\cdots,n-1$
$u_{n}$ $=1-\frac{2(y_{n}+1)}{\sum_{k}y_{k}^{2}+(y_{n}+1)^{2}}.$
Combining them together, we get the expression of the coordinates $x_{i}$’s in the upper half-space $\mathbb{H}^{n}$:
$x_{0}$ $=\frac{\sum_{k}y_{k}^{2}+y_{n}^{2}+1}{2y_{n}},$
$x_{k}$ $=\frac{y_{k}}{y_{n}},\qquad\qquad\qquad k=1,\cdots,n-1$
$x_{n}$ $=\frac{\sum_{k}y_{k}^{2}+y_{n}^{2}-1}{2y_{n}}.$
Proposition 4.12.
Consider the upper half-space model of $\mathbb{H}^{n}$. Let $\Sigma=S\cap\mathbb{H}^{n}$ be the intersection of $\mathbb{H}^{n}$ with a Euclidean $(n-1)$-sphere $S\subset\mathbb{R}^{n}$ and center in $\mathbb{H}^{n}$. Suppose $S$ makes an angle $\alpha$ with $\partial\mathbb{H}^{n}$. Then both the interior region and the exterior region with boundary $\Sigma$ are non-generic domains in $\mathbb{H}^{n}$. Moreover,
$\operatorname{ker}\Phi^{*}=\operatorname{span}\{x_{1},\cdots,x_{n-1},(2-\operatorname{cos}^{2}\alpha)x_{0}+\operatorname{cos}^{2}\alpha\,x_{n}\}.$
Here $(x_{0},x_{1},\cdots,x_{n})$ are the coordinates of $\mathbb{H}^{n}\subset\mathbb{R}^{1,n}$ in the hyperboloid model.
Proof.
We will just verify that the interior region bounded by $\Sigma$ is a non-generic domain in $\mathbb{H}^{n}$; the same calculation will give us that the exterior region bounded by $\Sigma$ is a non-generic domain as well.
Let us use $y_{i}$’s for the coordinates in $\mathbb{R}^{n}$. The metric of the upper half-space model of $\mathbb{H}^{n}$ is $g=y_{n}^{-2}\delta$, where $\delta$ is the Euclidean metric. Without loss of generality, we may assume the Euclidean radius of $S$ is 1, and that the center of $S$ is $O=(0,\cdots,0,\operatorname{cos}\alpha)$. Suppose $\vec{e}=\sum a_{i}\frac{\partial}{\partial y_{i}}$ is any vector with Euclidean norm 1. Then we may parametrize $\Sigma$ as $\{\vec{y}=O+\vec{e}\}$, which means the points on $\Sigma$ satisfy
$y_{k}$ $=a_{k}\qquad(k=1,\cdots,n-1)$
$y_{n}$ $=\operatorname{cos}\alpha+a_{n}$
Denote the outward unit normal at $\vec{y}$ as $\nu=c\vec{e}$, then
$1=g(\nu,\nu)=y_{n}^{-2}\delta(c\vec{e},c\vec{e})=c^{2}y_{n}^{-2}$
So
$\nu=y_{n}\vec{e}=(\operatorname{cos}\alpha+a_{n})\sum a_{i}\frac{\partial}{\partial y_{i}}$
Note that the mean curvature with respect to this normal is $H=(n-1)\operatorname{cos}\alpha$. Then
$(x_{0})_{\nu}-\frac{H}{n-1}x_{0}=y_{n}\sum a_{i}\frac{\partial}{\partial y_{i}}x_{0}-\operatorname{cos}\alpha\,x_{0}=-\frac{1}{2}\operatorname{cos}^{2}\alpha$
$(x_{k})_{\nu}-\frac{H}{n-1}x_{k}=y_{n}\sum a_{i}\frac{\partial}{\partial y_{i}}x_{k}-\operatorname{cos}\alpha\,x_{k}=a_{k}-y_{k}=0\quad(k=1,\cdots,n-1)$
$(x_{n})_{\nu}-\frac{H}{n-1}x_{n}=y_{n}\sum a_{i}\frac{\partial}{\partial y_{i}}x_{n}-\operatorname{cos}\alpha\,x_{n}=1-\frac{1}{2}\operatorname{cos}^{2}\alpha>0$
where we make use of the relation $\sum a_{i}^{2}=1$. Notice that the boundary equation is linear, and we have
$\left((2-\operatorname{cos}^{2}\alpha)x_{0}+\operatorname{cos}^{2}\alpha\,x_{n}\right)_{\nu}-\frac{H}{n-1}\left((2-\operatorname{cos}^{2}\alpha)x_{0}+\operatorname{cos}^{2}\alpha\,x_{n}\right)=0.$
So the region with boundary $\Sigma$ is a non-generic domain in $\mathbb{H}^{n}$, and
$\operatorname{ker}\Phi^{*}=\operatorname{span}\{x_{1},\cdots,x_{n-1},(2-\operatorname{cos}^{2}\alpha)x_{0}+\operatorname{cos}^{2}\alpha\,x_{n}\}.$
Remark 4.13.
Note the above proposition tells us that both the horospheres and the hyperspheres are boundaries of non-generic domains in $\mathbb{H}^{n}$.
Proposition 4.14.
Consider the upper half-space model of $\mathbb{H}^{n}$. Let $\Sigma=S\cap\mathbb{H}^{n}$ be the intersection of $\mathbb{H}^{n}$ with a Euclidean $(n-1)$-sphere $S\subset\mathbb{R}^{n}$ and center in $\mathbb{H}^{n}$. Suppose $S\subset\mathbb{H}^{n}$ is a geodesic sphere with Euclidean radius $1$ and the Euclidean center of $S$ relative to $\partial\mathbb{H}^{n}$ is at height $h$. Then both the interior region and the exterior region with boundary $\Sigma$ are non-generic domains in $\mathbb{H}^{n}$. Moreover,
$\operatorname{ker}\Phi^{*}=\operatorname{span}\{x_{1},\cdots,x_{n-1},(2-h^{2})x_{0}+h^{2}\,x_{n}\}.$
Here $(x_{0},x_{1},\cdots,x_{n})$ are the coordinates of $\mathbb{H}^{n}\subset\mathbb{R}^{1,n}$ in the hyperboloid model.
Proof.
Without loss of generality, we may assume the center of $S$ is $O=(0,\cdots,0,h)$. Then by replacing $\operatorname{cos}\alpha$ by $h$, the same calculation as above will give us that
$(x_{0})_{\nu}-\frac{H}{n-1}x_{0}=y_{n}\sum a_{i}\frac{\partial}{\partial y_{i}}x_{0}-hx_{0}=-\frac{1}{2}h^{2}<0$
$(x_{k})_{\nu}-\frac{H}{n-1}x_{k}=y_{n}\sum a_{i}\frac{\partial}{\partial y_{i}}x_{k}-hx_{k}=a_{k}-y_{k}=0\quad(k=1,\cdots,n-1)$
$(x_{n})_{\nu}-\frac{H}{n-1}x_{n}=y_{n}\sum a_{i}\frac{\partial}{\partial y_{i}}x_{n}-hx_{n}=1-\frac{1}{2}h^{2}$
The result follows immediately. ∎
Proposition 4.15.
Consider the upper half-space model of $\mathbb{H}^{n}$. Let $\Sigma=P\cap\mathbb{H}^{n}$ be the intersection of $\mathbb{H}^{n}$ with a Euclidean $(n-1)$-plane $P\subset\mathbb{R}^{n}$ that is parallel to $\partial\mathbb{H}^{n}$. Then the region with boundary $\Sigma$ is a non-generic domain in $\mathbb{H}^{n}$. Moreover,
$\operatorname{ker}\Phi^{*}=\operatorname{span}\{x_{1},\cdots,x_{n-1},x_{0}-x_{n}\}.$
Here $(x_{0},x_{1},\cdots,x_{n})$ are the coordinates of $\mathbb{H}^{n}\subset\mathbb{R}^{1,n}$ in the hyperboloid model.
Proof.
We will just verify that the lower region bounded by $\Sigma$ is a non-generic domain in $\mathbb{H}^{n}$; the same calculation will give us that the upper region bounded by $\Sigma$ is a non-generic domain as well.
We may parametrize $\Sigma$ as $\{y_{n}=c>0\}$, with the outward unit normal $\nu=y_{n}\frac{\partial}{\partial y_{n}}$. So the mean curvature with respect to this normal is
$H=y_{n}\left(H_{0}+(n-1)\left<\nabla(-\operatorname{ln}y_{n}),\frac{\partial}{\partial y_{n}}\right>_{\delta}\right)=-(n-1).$
Then
$(x_{0})_{\nu}-\frac{H}{n-1}x_{0}=y_{n}\frac{\partial}{\partial y_{n}}x_{0}+x_{0}=y_{n}>0$
$(x_{k})_{\nu}-\frac{H}{n-1}x_{k}=y_{n}\frac{\partial}{\partial y_{n}}x_{k}+x_{k}=-\frac{y_{k}}{y_{n}}+\frac{y_{k}}{y_{n}}=0\quad(k=1,\cdots,n-1)$
$(x_{n})_{\nu}-\frac{H}{n-1}x_{n}=y_{n}\frac{\partial}{\partial y_{n}}x_{n}+x_{n}=y_{n}>0$
The result follows immediately. ∎
Proposition 4.16.
Consider the upper half-space model of $\mathbb{H}^{n}$. Let $\Sigma=P\cap\mathbb{H}^{n}$ be the intersection of $\mathbb{H}^{n}$ with a Euclidean $(n-1)$-plane $P\subset\mathbb{R}^{n}$ that makes an angle $\alpha\in(0,\pi)$ with $\partial\mathbb{H}^{n}$. Then the region with boundary $\Sigma$ is a non-generic domain in $\mathbb{H}^{n}$. Moreover, if we parametrize $\Sigma$ as $\{\operatorname{cos}\alpha\,y_{n}=\operatorname{sin}\alpha\,y_{1}:y_{n}>0\}$, then
$\operatorname{ker}\Phi^{*}=\operatorname{span}\{x_{0},x_{2},\cdots,x_{n}\}.$
Here $(x_{0},x_{1},\cdots,x_{n})$ are the coordinates of $\mathbb{H}^{n}\subset\mathbb{R}^{1,n}$ in the hyperboloid model.
Proof.
We will just verify that the lower region bounded by $\Sigma$ is a non-generic domain in $\mathbb{H}^{n}$; the same calculation will give us that the upper region bounded by $\Sigma$ is a non-generic domain as well.
Without loss of generality, we may parametrize $\Sigma$ as $\{\operatorname{cos}\alpha\,y_{n}=\operatorname{sin}\alpha\,y_{1}:y_{n}>0\}$, with the outward unit normal $\nu=y_{n}(-\operatorname{sin}\alpha\frac{\partial}{\partial y_{1}}+\operatorname{cos}\alpha\frac{\partial}{\partial y_{n}})$. So the mean curvature with respect to this normal is
$H=y_{n}\left(H_{0}+(n-1)\left<\nabla(-\operatorname{ln}y_{n}),-\operatorname{sin}\alpha\frac{\partial}{\partial y_{1}}+\operatorname{cos}\alpha\frac{\partial}{\partial y_{n}}\right>_{\delta}\right)=-(n-1)\operatorname{cos}\alpha.$
Then
$(x_{0})_{\nu}-\frac{H}{n-1}x_{0}=y_{n}(-\operatorname{sin}\alpha\frac{\partial}{\partial y_{1}}+\operatorname{cos}\alpha\frac{\partial}{\partial y_{n}})x_{0}+\operatorname{cos}\alpha\,x_{0}=0$
$(x_{1})_{\nu}-\frac{H}{n-1}x_{1}=y_{n}(-\operatorname{sin}\alpha\frac{\partial}{\partial y_{1}}+\operatorname{cos}\alpha\frac{\partial}{\partial y_{n}})x_{1}+\operatorname{cos}\alpha\,x_{1}=-\operatorname{sin}\alpha<0$
$(x_{k})_{\nu}-\frac{H}{n-1}x_{k}=y_{n}(-\operatorname{sin}\alpha\frac{\partial}{\partial y_{1}}+\operatorname{cos}\alpha\frac{\partial}{\partial y_{n}})x_{k}+\operatorname{cos}\alpha\,x_{k}=0\,\,(k=2,\cdots,n-1)$
$(x_{n})_{\nu}-\frac{H}{n-1}x_{n}=y_{n}(-\operatorname{sin}\alpha\frac{\partial}{\partial y_{1}}+\operatorname{cos}\alpha\frac{\partial}{\partial y_{n}})x_{n}+\operatorname{cos}\alpha\,x_{n}=0$
The result follows immediately. ∎
As a summary, we have the following classification result:
Theorem 4.17 (Classification of simple non-generic domains in $\mathbb{H}^{n}$).
There are four types of boundaries of non-generic domains in $\mathbb{H}^{n}$:
(i) the horosphere;       (ii) the hypersphere;
(iii) the geodesic sphere;  (iv) the intersection with $\mathbb{H}^{n}$ of hyperplanes of $\mathbb{R}^{n}$.
We notice Hijazi-Montiel-Raulot [H-M-R] had the following rigidity result for asymptotically hyperbolic manifolds with inner boundary:
Proposition 4.18 (Hijazi-Montiel-Raulot).
Let $(M^{3},g)$ be a three-dimensional complete AH manifold with scalar curvature satisfying $R_{g}\geq-6$ and compact inner boundary $\Sigma$ homeomorphic to a $2$-sphere whose mean curvature is such that
$H\leq 2\sqrt{\frac{4\pi}{\operatorname{Area}_{g}(\Sigma)}+1}.$
Then the energy-momentum vector $\mathbf{p}_{g}$ is time-like future-directed or zero. Moreover, if it vanishes then $(M^{3},g)$ is isometric to the complement of a geodesic ball in $\mathbb{H}^{3}$.
On the other hand, Almaraz-de Lima [A-dL] had the following rigidity result for asymptotically hyperbolic manifolds with a non-compact boundary:
Proposition 4.19 (Almaraz-de Lima).
Let $(M,g,\Sigma)$ be an asymptotically hyperbolic spin manifold with boundary, with $R_{g}\geq-n(n-1)$ and $H_{g}\geq 0$. Assume further that $g$ agrees with the reference hyperbolic metric $b$ in a neighborhood of infinity. Then $(M,g,\Sigma)=(\mathbb{H}^{n}_{+},b,\partial\mathbb{H}^{n}_{+})$ isometrically.
These results together give us a very good explanation of non-generic domains in $\mathbb{H}^{n}$.
4.4. The Schwarzschild manifold
Finally, we would like to determine the non-generic domains in the Schwarzschild manifold. We have already known that the umbilic hypersurfaces in $\mathbb{R}^{n}$ are spheres and hyperplanes, and that umbilic hypersurfaces remain umbilic after the conformal change [dC]. Since the $n$-dimensional Schwarzschild manifold is conformal to $\mathbb{R}^{n}\backslash\{\mathbf{0}\}$, we get the umbilic hypersurfaces in the Schwarzschild manifold are Euclidean spheres and hyperplanes.
Proposition 4.20.
Consider the Schwarzschild metric $g^{S}=\left(1+\frac{m}{2r^{n-2}}\right)^{\frac{4}{n-2}}\delta$ on $\mathbb{R}^{n}\backslash\{\mathbf{0}\}$, where $r=|\vec{x}|$. Denote $\Sigma$ as the Euclidean sphere centered at $\mathbf{0}$ with radius $r_{\pm}=\left(\frac{(n-1)\pm\sqrt{n^{2}-2n}}{2}m\right)^{\frac{1}{n-2}}$. Then the region with boundary $\Sigma$ is a non-generic domain in the Schwarzschild manifold. Moreover,
$\operatorname{ker}\Phi^{*}=\operatorname{span}\left\{\frac{1-\frac{m}{2r^{n-2}}}{1+\frac{m}{2r^{n-2}}}\right\}.$
Proof.
As mentioned in [C], the function $u_{0}=\frac{1-\frac{m}{2r}}{1+\frac{m}{2r}}\in\operatorname{ker}L_{g^{S}}^{*}$ is a possible static potential in $3$-dimensional Schwarzschild manifold. Analogously, the function $u=\frac{1-\frac{m}{2r^{n-2}}}{1+\frac{m}{2r^{n-2}}}\in\operatorname{ker}L_{g^{S}}^{*}$ is a possible static potential in $n$-dimensional Schwarzschild manifold. (This can also be found using the method introduced in [B-M].) Indeed, it is the only possible static potential in connected open domains, as found by Corvino [C-E-M]. Now let us verify whether the function satisfies the boundary condition on Euclidean spheres centered at $\mathbf{0}$.
Suppose $\vec{e}$ is any vector with Euclidean norm 1. Then we may parametrize the Euclidean sphere as $\{\vec{x}=r\vec{e}\}$. Denote the outward unit normal at $\vec{x}$ as $\nu=c\vec{e}$, then
$1=g^{S}(\nu,\nu)=\left(1+\frac{m}{2r^{n-2}}\right)^{\frac{4}{n-2}}\delta(c\vec{e},c\vec{e})=c^{2}\left(1+\frac{m}{2r^{n-2}}\right)^{\frac{4}{n-2}}$
So
$\nu=\left(1+\frac{m}{2r^{n-2}}\right)^{-\frac{2}{n-2}}\vec{e}$
Note that the mean curvature with respect to this normal is
$H$ $=\left(1+\frac{m}{2r^{n-2}}\right)^{-\frac{2}{n-2}}\left(H_{0}+(n-1)\left<\nabla\left(\frac{2}{n-2}\operatorname{ln}(1+\frac{m}{2r^{n-2}})\right),\vec{e}\right>_{\delta}\right)$
$=\left(1+\frac{m}{2r^{n-2}}\right)^{-\frac{2}{n-2}}\left(\frac{n-1}{r}-(n-1)\left<\frac{m\vec{x}}{r^{n}(1+\frac{m}{2r^{n-2}})},\vec{e}\right>_{\delta}\right)$
$=\left(1+\frac{m}{2r^{n-2}}\right)^{-\frac{2}{n-2}}\left(\frac{n-1}{r}-\frac{(n-1)m}{r^{n-1}(1+\frac{m}{2r^{n-2}})}\right)$
$=(n-1)\left(1+\frac{m}{2r^{n-2}}\right)^{-\frac{2}{n-2}}\frac{1-\frac{m}{2r^{n-2}}}{r(1+\frac{m}{2r^{n-2}})}$
(4.1) $=(n-1)\left(1+\frac{m}{2r^{n-2}}\right)^{-\frac{2}{n-2}}r^{-1}u$
Then
$u_{\nu}-\frac{H}{n-1}u$ $=\left(1+\frac{m}{2r^{n-2}}\right)^{-\frac{2}{n-2}}\vec{e}\left(\frac{1-\frac{m}{2r^{n-2}}}{1+\frac{m}{2r^{n-2}}}\right)-\left(1+\frac{m}{2r^{n-2}}\right)^{-\frac{2}{n-2}}r^{-1}u^{2}$
$=\left(1+\frac{m}{2r^{n-2}}\right)^{-\frac{2}{n-2}}\left<\frac{(n-2)m\vec{x}}{r^{n}(1+\frac{m}{2r^{n-2}})^{2}},\vec{e}\right>_{\delta}-\frac{(1-\frac{m}{2r^{n-2}})^{2}}{r(1+\frac{m}{2r^{n-2}})^{2+\frac{2}{n-2}}}$
$=\frac{(n-2)m}{r^{n-1}(1+\frac{m}{2r^{n-2}})^{2+\frac{2}{n-2}}}-\frac{\frac{m^{2}}{4r^{n-2}}-m+r^{n-2}}{r^{n-1}(1+\frac{m}{2r^{n-2}})^{2+\frac{2}{n-2}}}$
$=\frac{-\frac{m^{2}}{4r^{n-2}}+(n-1)m-r^{n-2}}{r^{n-1}(1+\frac{m}{2r^{n-2}})^{2+\frac{2}{n-2}}}$
Thus $u_{\nu}-\frac{H}{n-1}u=0$ if and only if
$r^{n-2}=\frac{(n-1)\pm\sqrt{n^{2}-2n}}{2}m.$
This means the region with boundary $\Sigma$ is a non-generic domain in $(\mathbb{R}^{n}\backslash\{\mathbf{0}\},g^{S})$. ∎
As a non-generic domain, one would expect it to have some special properties in the Schwarzschild manifold. Unfortunately, neither $r_{\pm}$ in Proposition 4.20 would give us the horizon of the Schwarzschild manifold, as the mean curvature $H$ cannot be 00 (otherwise we would get $u\equiv 0$ from (4.4)). However, in the next proposition we will see that the two radii $r_{\pm}$ do have special geometric and physical meanings.
Proposition 4.21.
The above two radii $r_{\pm}$ are the critical points of the mean curvature $H=H(r)$ on the Euclidean sphere (centered at $\mathbf{0}$) with radius $r$.
Proof.
As we can see from (4.4), the mean curvature on the Euclidean sphere (centered at $\mathbf{0}$) with radius $r$ is a function depending only on $r$, that is, $H=H(r)$.
Let us calculate its derivative $H^{\prime}$, and denote $\alpha=\frac{m}{2r^{n-2}}$.
$H^{\prime}$ $=mr^{1-n}(1+\alpha)^{-1}H-r^{-1}H+\frac{(n-2)mr^{1-n}}{(1-\alpha)(1+\alpha)}H$
$=H\frac{1}{r(1-\alpha)(1+\alpha)}\left[2\alpha(1-\alpha)-(1-\alpha)(1+\alpha)+(n-2)2\alpha\right]$
$=H\frac{1}{r(1-\alpha)(1+\alpha)}\left(-\alpha^{2}+2(n-1)\alpha-1\right).$
Note that $H\neq 0$, so $H^{\prime}=0$ if and only if
$\alpha^{2}-2(n-1)\alpha+1=0,$
that is,
$\alpha=\frac{m}{2r^{n-2}}=(n-1)\pm\sqrt{n^{2}-2n}.$
This means, $H^{\prime}=0$ if and only if $r=r_{\pm}$. ∎
From (4.4) we can see, when $r\rightarrow 0$ or $r\rightarrow\infty$, $H\rightarrow 0$. And $H=0$ on the horizon ($r^{n-2}=\frac{m}{2}$). So if the mass $m>0$,
(i) $r_{-}$ (the one inside the horizon) gives us the minimal mean curvature $H_{\text{min}}$;
(ii) $r_{+}$ (the one outside the horizon) gives us the maximal mean curvature $H_{\text{max}}$.
Moreover, $(r_{-}\cdot r_{+})^{n-2}=\frac{m^{2}}{4}$; in other words, the two spheres are symmetric w.r.t. the horizon.
$r_{-}$$r_{+}$$\mathbf{0}$HorizonPhoton sphere
Figure 2. The Schwarzschild photon sphere and the horizon
In 3D, physicists call the one outside the horizon with radius $r_{+}=\frac{2+\sqrt{3}}{2}m$ the Schwarzschild photon sphere [Claudel-V-E, Cederbaum]. Roughly speaking, it behaves like a borderline in the following sense:
(i) Any endless geodesic starting at some point outside the photon sphere and initially directed outwards will continue outwards and escape to infinity.
(ii) Any endless geodesic starting at some point between the horizon and the photon sphere and initially directed inwards will continue inwards and fall into the black hole.
(iii) Any geodesic starting at some point on the photon sphere and initially tangent to the photon sphere will remain in the photon sphere.
The Schwarzschild photon sphere thus models (an embedded submanifold ruled by) photons spiraling around the central black hole “at a fixed distance”. The Schwarzschild photon sphere and the notion of trapped null geodesics in general are crucially relevant for questions of dynamical stability in the context of the Einstein equations. Moreover, photon spheres are related to the existence of relativistic images in the context of gravitational lensing. On the other hand, Cederbaum and Galloway [Cederbaum, Cederbaum-G] also found rigidity results on the exterior region of the photon sphere, which happens to be a non-generic domain in our setting.
Remark 4.22.
Even though the Euclidean spheres not centered at $\mathbf{0}$ are umbilic hypersurfaces in the Schwarzschild manifold, they are not boundaries of non-generic domains. This is mainly because the function $r$ is no longer a constant on the sphere, thus $H$ is not constant. More precisely, we may assume the sphere is centered at $\vec{e}$ ($\vec{e}$ has Euclidean norm $1$) with radius $\rho$. Then let us look at two different points on the sphere, namely $x_{1}=\vec{e}+\rho\vec{e}$ and $x_{2}=\vec{e}-\rho\vec{e}$. Following the same calculation as above, we will see the two $\rho$’s satisfying the boundary condition are different.
Proposition 4.23.
Denote $\Sigma$ as the Euclidean hyperplane containing $\mathbf{0}$. Then the region with boundary $\Sigma$ is a non-generic domain in the Schwarzschild manifold. Moreover,
$\operatorname{ker}\Phi^{*}=\operatorname{span}\left\{\frac{1-\frac{m}{2r^{n-2}}}{1+\frac{m}{2r^{n-2}}}\right\}.$
Proof.
First note that $\Sigma$ is complete and has two ends in the Schwarzschild manifold. Now let us verify whether the static potential mentioned above satisfies the boundary condition.
Without loss of generality, we may parametrize $\Sigma$ as $\{x_{n}=c\}$. Then the outward unit normal is
$\nu=\left(1+\frac{m}{2r^{n-2}}\right)^{-\frac{2}{n-2}}\frac{\partial}{\partial x_{n}}$
and the mean curvature is
$H$ $=\left(1+\frac{m}{2r^{n-2}}\right)^{-\frac{2}{n-2}}\left(H_{0}+(n-1)\left<\nabla\left(\frac{2}{n-2}\operatorname{ln}(1+\frac{m}{2r^{n-2}})\right),\frac{\partial}{\partial x_{n}}\right>_{\delta}\right)$
$=-\left(1+\frac{m}{2r^{n-2}}\right)^{-\frac{2}{n-2}}\left((n-1)\frac{mx_{n}}{r^{n}(1+\frac{m}{2r^{n-2}})}\right)$
$=-(n-1)\frac{mx_{n}}{r^{n}(1+\frac{m}{2r^{n-2}})^{1+\frac{2}{n-2}}}$
Then
$u_{\nu}-\frac{H}{n-1}u$ $=\left(1+\frac{m}{2r^{n-2}}\right)^{-\frac{2}{n-2}}\frac{(n-2)mx_{n}}{r^{n}(1+\frac{m}{2r^{n-2}})^{2}}+\frac{mx_{n}}{r^{n}(1+\frac{m}{2r^{n-2}})^{1+\frac{2}{n-2}}}\left(\frac{1-\frac{m}{2r^{n-2}}}{1+\frac{m}{2r^{n-2}}}\right)$
$=\frac{(n-2)mx_{n}}{r^{n}(1+\frac{m}{2r^{n-2}})^{2+\frac{2}{n-2}}}+\frac{mx_{n}(1-\frac{m}{2r^{n-2}})}{r^{n}(1+\frac{m}{2r^{n-2}})^{2+\frac{2}{n-2}}}$
$=\frac{mc((n-1)-\frac{m}{2r^{n-2}})}{r^{n}(1+\frac{m}{2r^{n-2}})^{2+\frac{2}{n-2}}}$
Thus $u_{\nu}-\frac{H}{n-1}u=0$ if and only if $c=0$. This means the region with boundary $\Sigma$ is a non-generic domain in $(\mathbb{R}^{n}\backslash\{\mathbf{0}\},g^{S})$. ∎
As a summary, we have the following classification result:
Theorem 4.24 (Classification of simple non-generic domains in the Schwarzschild manifold).
There are two types of boundaries of non-generic domains in the Schwarzschild manifold:
(i) the Euclidean spheres centered at $\mathbf{0}$ with radii $r_{\pm}$;
(ii) the Euclidean hyperplanes containing $\mathbf{0}$.
4.5. Non-generic domains with multiple boundary components
For the above examples in space forms and in the Schwarzschild manifold, we are just considering simple non-generic domains with only one boundary component. If a non-generic domain has multiple boundary components, then it can be viewed as the intersection of simple non-generic domains, as long as their boundaries do not intersect with each other. In this case, the space of static potentials is the intersection of those corresponding to each simple non-generic domain. But of course we will need to make sure the eventual space of static potentials has dimension at least $1$.
Example 4.25.
Consider the standard metric $g_{\mathbb{S}^{n}}$ on the unit sphere $\mathbb{S}^{n}$, and let $(x_{1},\cdots,x_{n+1})$ denote the coordinates of $\mathbb{S}^{n}\subset\mathbb{R}^{n+1}$. For $\epsilon>0$ small, define $\Sigma_{j}\subset\mathbb{S}^{n}$ by
$\Sigma_{j}=\{\mathbf{x}\in\mathbb{S}^{n}:x_{j}=1-\epsilon\}\qquad(j=1,\cdots,n)$
Then $\Sigma=\bigcup\limits_{j=1}^{n}\Sigma_{j}$ is a hypersurface with $n$ components. Denote the connected domain with boundary $\Sigma$ as $\Omega$. Then by Theorem 4.3 we know $(\Omega,\Sigma)$ is a non-generic domain in $\mathbb{S}^{n}$, and moreover
$\operatorname{ker}\Phi^{*}=\operatorname{span}\{x_{n+1}\}.$
Acknowledgments
I would like to thank my PhD advisor Rick Schoen for his wisdom and patience during the preparation of my dissertation.